Demand Curve: p=95-q Supply Curve: p=15+3q What is the equilibrium price and quantity?
Answer #1

At the equilibrium the demand and supply are equal .

95 - q = 15 + 3q.

= 95-15 = 3q+q

80 = 4q

Q = 80/4

Q = 20.

Price = 95 -20

Price = 75.

At the equilibrium the price will be 75 and quantity will be 20.